# FedP2EFT: Federated Learning to Personalize Parameter Efficient Fine-Tuning for Multilingual LLMs

### v1

@@ Appendix B Additional Results
$\mathbf{r}$ Approach bg hi es el vi tr de ur en zh th sw ar fr ru Wins
2 LoRA 49.33±0.09 45.93±0.25 51.13±0.09 48.53±0.25 48.53±0.25 45.27±0.25 49.80±0.16 46.87±0.34 55.40±0.28 49.33±0.25 44.40±0.00 41.20±0.16 46.87±0.09 48.33±0.09 45.13±0.09 0
AdaLoRA 49.33±0.09 45.27±0.09 50.27±0.19 47.80±0.16 47.27±0.09 44.60±0.00 49.33±0.34 46.60±0.16 54.87±0.25 48.73±0.19 44.73±0.09 40.07±0.09 46.67±0.09 48.67±0.09 44.53±0.09 0
BayesTune-LoRA 49.13±0.09 45.13±0.19 50.20±0.00 48.00±0.16 47.07±0.19 44.67±0.09 49.07±0.25 46.47±0.19 54.67±0.09 48.73±0.19 44.80±0.16 40.07±0.09 46.60±0.16 48.73±0.09 44.27±0.09 0
FedL2P 54.40±0.43 49.93±0.25 56.80±0.33 51.40±0.33 54.00±0.71 48.27±0.50 54.00±0.59 50.13±0.41 59.93±0.41 52.47±0.25 48.00±0.28 44.53±0.34 50.33±0.41 53.00±0.28 49.13±0.74 6
FedP2EFT 54.00±0.59 51.13±0.34 57.67±0.09 51.13±0.68 54.53±0.25 49.73±0.09 53.33±0.66 53.13±0.98 59.60±0.33 51.53±0.38 48.87±0.19 46.53±0.66 49.00±0.43 56.33±0.68 52.47±0.41 9
4 LoRA 49.87±0.09 47.07±0.09 52.27±0.34 49.67±0.19 49.87±0.25 46.47±0.34 50.33±0.09 48.20±0.16 56.53±0.62 49.80±0.16 45.00±0.43 41.40±0.00 47.93±0.09 48.73±0.09 45.80±0.16 0
AdaLoRA 49.27±0.09 45.20±0.16 50.07±0.25 47.93±0.25 47.27±0.09 44.60±0.00 49.27±0.19 46.53±0.09 54.73±0.25 48.60±0.16 44.60±0.16 40.07±0.19 46.47±0.09 48.80±0.00 44.53±0.34 0
BayesTune-LoRA 49.33±0.19 45.13±0.09 50.00±0.16 48.20±0.16 47.13±0.09 44.67±0.09 49.40±0.00 46.40±0.16 54.93±0.09 48.87±0.09 44.60±0.00 40.13±0.09 46.60±0.00 48.53±0.09 44.40±0.00 0
FedL2P 54.73±0.38 50.93±0.66 57.27±0.34 51.47±0.34 54.67±1.09 47.67±0.25 54.20±0.43 50.93±0.25 59.80±0.59 52.73±0.09 47.87±0.50 44.80±0.00 50.20±0.28 53.47±0.50 49.53±0.93 6
FedP2EFT 55.33±0.34 50.60±0.91 58.20±0.16 51.33±0.41 54.60±0.28 48.80±0.28 52.73±0.57 53.47±0.47 59.13±0.68 51.53±0.34 48.47±0.41 47.20±0.59 49.00±0.16 56.73±0.66 52.13±0.41 9
8 LoRA 51.00±0.33 48.13±0.50 54.00±0.33 50.87±0.34 51.20±0.28 46.93±0.25 52.13±1.05 48.27±0.25 58.33±0.19 50.13±0.19 45.80±0.16 42.40±0.00 48.13±1.05 50.93±0.09 46.80±0.59 0
AdaLoRA 49.20±0.16 45.40±0.16 50.13±0.19 48.13±0.09 47.07±0.09 44.73±0.09 49.13±0.09 46.33±0.09 54.93±0.19 48.47±0.34 44.60±0.00 40.13±0.09 46.67±0.09 48.67±0.19 44.33±0.25 0
BayesTune-LoRA 49.27±0.09 45.07±0.19 50.27±0.09 47.93±0.09 47.20±0.16 44.73±0.09 49.27±0.19 46.33±0.19 54.93±0.09 48.53±0.09 44.53±0.09 40.33±0.09 46.67±0.09 48.53±0.09 44.53±0.38 0
FedL2P 54.13±0.25 49.60±1.14 57.47±0.82 50.87±0.50 53.20±0.28 47.60±0.43 54.20±0.16 49.87±0.09 59.53±0.62 52.27±0.90 47.47±0.25 43.60±0.71 50.27±0.52 52.60±0.57 48.80±0.57 5
FedP2EFT 53.40±0.00 50.60±0.33 57.67±0.62 50.53±0.34 55.13±0.38 48.00±0.75 52.53±0.34 51.87±0.19 57.00±1.02 50.67±0.09 47.60±0.00 46.67±0.52 48.47±0.66 55.00±0.33 50.73±0.75 9
16 LoRA 52.73±0.25 48.80±0.00 55.47±0.25 51.13±1.05 53.53±0.41 48.87±0.75 54.40±0.91 49.40±0.99 58.87±1.11 51.80±0.49 46.93±0.09 43.60±0.99 49.13±1.05 52.13±0.81 48.93±2.03 3
AdaLoRA 49.13±0.09 45.20±0.00 50.20±0.16 47.93±0.25 46.93±0.19 44.53±0.09 49.33±0.19 46.53±0.09 54.80±0.16 48.73±0.09 44.73±0.19 40.13±0.09 46.73±0.25 48.67±0.09 44.40±0.28 0
BayesTune-LoRA 49.33±0.25 45.33±0.19 50.20±0.00 47.93±0.09 47.53±0.25 44.80±0.16 49.40±0.33 46.53±0.09 55.27±0.09 48.87±0.34 44.47±0.09 40.20±0.28 46.53±0.09 48.73±0.25 44.73±0.25 0
FedL2P 52.53±1.16 49.00±0.28 56.40±1.40 50.47±0.96 52.47±0.68 47.80±0.28 53.93±0.38 48.80±0.85 59.07±0.41 51.73±0.75 47.13±0.82 43.00±1.14 49.27±0.84 52.07±1.16 48.33±0.74 3
FedP2EFT 53.73±0.52 49.87±0.50 57.07±0.52 49.87±0.25 54.67±0.90 47.67±0.41 52.27±0.38 52.27±0.52 57.87±0.68 50.33±0.66 47.60±0.33 46.20±0.16 49.00±0.49 54.60±0.71 49.67±0.57 9
@@
Table 10: Average METEOR/ROUGE-1/ROUGE-L of each language for seen clients of our Fed-Aya setup. The pretrained MobileLLaMA-1.4B model is trained using Standard FL with LoRA following FedLLM-Bench (Ye et al., 2024) and the resulting base model is personalized to each client given a baseline approach.
$\mathbf{r}$ Approach te ar es en fr zh pt Wins
2 LoRA 0.1207/0.0545/0.0524 0.1835/0.0210/0.0205 0.3228/0.3519/0.3251 0.2457/0.3121/0.2525 0.2049/0.2484/0.2002 0.0616/0.0874/0.0873 0.2631/0.3266/0.2922 1
AdaLoRA 0.1238/0.0607/0.0586 0.1819/0.0223/0.0218 0.3288/0.3573/0.3315 0.2459/0.3172/0.2524 0.1963/0.2422/0.1915 0.0689/0.0832/0.0832 0.2745/0.3327/0.2986 0
BayesTune-LoRA 0.1245/0.0605/0.0580 0.1813/0.0181/0.0178 0.2941/0.3317/0.3063 0.2345/0.2892/0.2367 0.1885/0.2430/0.1970 0.0643/0.0805/0.0805 0.2408/0.2971/0.2645 0
FedL2P 0.1451/0.0747/0.0725 0.2017/0.0219/0.0213 0.3321/0.3523/0.3245 0.2635/0.3307/0.2692 0.2298/0.2467/0.2034 0.0544/0.0803/0.0803 0.2780/0.3133/0.2832 1
FedP2EFT 0.1266/0.0629/0.0606 0.2081/0.0254/0.0248 0.3425/0.3663/0.3376 0.2745/0.3469/0.2831 0.2342/0.2766/0.2246 0.0524/0.0777/0.0777 0.2846/0.3403/0.3066 5
4 LoRA 0.1232/0.0570/0.0537 0.1861/0.0202/0.0197 0.3284/0.3555/0.3291 0.2541/0.3202/0.2585 0.2037/0.2475/0.1990 0.0559/0.0886/0.0886 0.2734/0.3314/0.2975 1
AdaLoRA 0.1240/0.0596/0.0574 0.1858/0.0180/0.0177 0.3310/0.3548/0.3287 0.2448/0.3111/0.2500 0.1892/0.2331/0.1859 0.0617/0.0852/0.0851 0.2640/0.3250/0.2934 0
BayesTune-LoRA 0.1214/0.0548/0.0532 0.1912/0.0201/0.0195 0.3042/0.3405/0.3150 0.2405/0.3022/0.2440 0.1973/0.2393/0.1925 0.0670/0.0806/0.0806 0.2468/0.3057/0.2737 0
FedL2P 0.1258/0.0616/0.0592 0.1805/0.0217/0.0208 0.3260/0.3568/0.3298 0.2519/0.3108/0.2498 0.2037/0.2493/0.2027 0.0484/0.0798/0.0798 0.2626/0.3174/0.2836 0
FedP2EFT 0.1350/0.0722/0.0692 0.2218/0.0275/0.0269 0.3448/0.3712/0.3419 0.2796/0.3516/0.2883 0.2469/0.2753/0.2244 0.0554/0.0843/0.0843 0.2911/0.3397/0.3060 6
8 LoRA 0.1241/0.0522/0.0493 0.2059/0.0189/0.0184 0.3435/0.3688/0.3427 0.2686/0.3400/0.2771 0.2197/0.2556/0.2037 0.0583/0.0886/0.0884 0.2822/0.3418/0.3076 0
AdaLoRA 0.1245/0.0623/0.0599 0.1771/0.0189/0.0184 0.3215/0.3485/0.3225 0.2461/0.3101/0.2512 0.1799/0.2287/0.1839 0.0613/0.0806/0.0806 0.2607/0.3194/0.2879 0
BayesTune-LoRA 0.1246/0.0591/0.0572 0.2046/0.0192/0.0189 0.3247/0.3520/0.3284 0.2430/0.3095/0.2496 0.2132/0.2542/0.2059 0.0601/0.0818/0.0818 0.2588/0.3171/0.2871 0
FedL2P 0.1316/0.0687/0.0661 0.1855/0.0218/0.0215 0.3272/0.3535/0.3280 0.2696/0.3304/0.2711 0.2109/0.2558/0.2059 0.0510/0.0816/0.0816 0.2750/0.3252/0.2897 1
FedP2EFT 0.1327/0.0662/0.0644 0.2304/0.0253/0.0246 0.3474/0.3847/0.3531 0.2941/0.3656/0.2996 0.2553/0.2829/0.2268 0.0538/0.0814/0.0814 0.2945/0.3452/0.3108 5
16 LoRA 0.1217/0.0562/0.0536 0.2080/0.0221/0.0218 0.3387/0.3616/0.3352 0.2757/0.3431/0.2807 0.2497/0.2880/0.2337 0.0553/0.0844/0.0844 0.2902/0.3449/0.3106 2
AdaLoRA 0.1251/0.0624/0.0602 0.1676/0.0198/0.0192 0.3048/0.3329/0.3037 0.2391/0.2985/0.2416 0.1821/0.2309/0.1866 0.0575/0.0815/0.0815 0.2530/0.3099/0.2784 0
BayesTune-LoRA 0.1374/0.0745/0.0720 0.2119/0.0175/0.0172 0.3358/0.3649/0.3397 0.2587/0.3189/0.2577 0.2222/0.2603/0.2113 0.0520/0.0824/0.0824 0.2862/0.3450/0.3109 0
FedL2P 0.1559/0.0827/0.0799 0.2013/0.0228/0.0226 0.3278/0.3541/0.3268 0.2772/0.3278/0.2693 0.2306/0.2346/0.1925 0.0506/0.0838/0.0838 0.2817/0.3179/0.2851 1
FedP2EFT 0.1309/0.0663/0.0638 0.2359/0.0258/0.0252 0.3447/0.3778/0.3463 0.2802/0.3485/0.2858 0.2473/0.2775/0.2208 0.0538/0.0835/0.0835 0.2975/0.3491/0.3157 4
@@
Table 11: Average METEOR/ROUGE-1/ROUGE-L of each language for seen clients of our Fed-Aya setup. The pretrained MobileLLaMA-1.4B model is trained using Standard FL with LoRA following FedLLM-Bench (Ye et al., 2024) and the resulting base model is personalized to each client given a baseline approach.
$\mathbf{r}$ Approach te ar es en fr zh pt ru Wins
2 LoRA 0.0531/0.0042/0.0042 0.1032/0.0524/0.0524 0.3547/0.4490/0.4020 0.2385/0.2713/0.2356 0.0490/0.0000/0.0000 0.1381/0.0088/0.0088 0.1865/0.2184/0.2070 0.0921/0.0556/0.0556 1
AdaLoRA 0.0312/0.0000/0.0000 0.1000/0.0520/0.0520 0.3244/0.4250/0.3821 0.2165/0.2606/0.2197 0.0490/0.0513/0.0513 0.1609/0.0089/0.0089 0.1904/0.2258/0.2129 0.0992/0.0556/0.0556 1
BayesTune-LoRA 0.0276/0.0000/0.0000 0.0934/0.0501/0.0476 0.3560/0.4084/0.3716 0.1786/0.2024/0.1688 0.0450/0.0000/0.0000 0.1566/0.0109/0.0092 0.1469/0.1628/0.1500 0.0873/0.0556/0.0556 0
FedL2P 0.1199/0.0042/0.0042 0.1399/0.0206/0.0206 0.3923/0.4587/0.4194 0.2688/0.3049/0.2650 0.0024/0.0011/0.0011 0.0888/0.0121/0.0121 0.2022/0.2170/0.2063 0.0850/0.0556/0.0556 2
FedP2EFT 0.1105/0.0132/0.0114 0.1127/0.0483/0.0483 0.3812/0.4546/0.4209 0.3047/0.3354/0.3006 0.0490/0.0000/0.0000 0.0650/0.0069/0.0069 0.1997/0.2505/0.2384 0.0956/0.0556/0.0556 3
4 LoRA 0.0687/0.0075/0.0075 0.0989/0.0581/0.0556 0.3244/0.4086/0.3749 0.2297/0.2726/0.2370 0.0490/0.0513/0.0513 0.1530/0.0089/0.0089 0.1872/0.2267/0.2124 0.0974/0.0556/0.0556 0
AdaLoRA 0.0263/0.0000/0.0000 0.0894/0.0634/0.0610 0.3336/0.4322/0.3883 0.1762/0.2105/0.1765 0.0490/0.0000/0.0000 0.1630/0.0066/0.0049 0.1832/0.2096/0.1980 0.1156/0.0556/0.0556 0
BayesTune-LoRA 0.0471/0.0075/0.0075 0.0938/0.0544/0.0544 0.3357/0.4058/0.3629 0.1657/0.1964/0.1647 0.1155/0.0635/0.0635 0.1557/0.0115/0.0099 0.1511/0.1636/0.1517 0.0858/0.0222/0.0222 1
FedL2P 0.0569/0.0075/0.0075 0.1002/0.0610/0.0610 0.3098/0.4066/0.3837 0.2494/0.2901/0.2569 0.1157/0.0741/0.0741 0.1585/0.0041/0.0041 0.1724/0.1920/0.1806 0.0804/0.0556/0.0556 1
FedP2EFT 0.1121/0.0212/0.0212 0.1537/0.0333/0.0333 0.3725/0.4622/0.4273 0.2945/0.3276/0.2821 0.0490/0.0000/0.0000 0.0722/0.0069/0.0069 0.1903/0.2547/0.2438 0.1258/0.0556/0.0556 5
8 LoRA 0.1022/0.0042/0.0042 0.1206/0.0549/0.0549 0.3401/0.4382/0.3977 0.2722/0.3127/0.2726 0.0490/0.0000/0.0000 0.1241/0.0089/0.0089 0.1879/0.2331/0.2202 0.0761/0.0556/0.0556 1
AdaLoRA 0.0218/0.0000/0.0000 0.0897/0.0364/0.0364 0.3383/0.4295/0.3920 0.1786/0.2150/0.1795 0.1221/0.0833/0.0833 0.1470/0.0066/0.0049 0.1773/0.1909/0.1770 0.1124/0.0556/0.0556 1
BayesTune-LoRA 0.0529/0.0042/0.0042 0.0970/0.0557/0.0557 0.3544/0.4257/0.3789 0.1891/0.2235/0.1860 0.1262/0.0784/0.0784 0.1522/0.0068/0.0068 0.1365/0.1603/0.1483 0.0895/0.0556/0.0556 1
FedL2P 0.0934/0.0042/0.0042 0.1138/0.0495/0.0495 0.3237/0.4366/0.4068 0.2438/0.2849/0.2451 0.0490/0.0000/0.0000 0.1408/0.0139/0.0139 0.1946/0.2280/0.2130 0.1187/0.0556/0.0556 1
FedP2EFT 0.1099/0.0132/0.0114 0.1565/0.0082/0.0082 0.4471/0.5240/0.4876 0.2798/0.2851/0.2475 0.0490/0.0000/0.0000 0.1071/0.0069/0.0069 0.2160/0.2700/0.2559 0.0862/0.0465/0.0465 3
16 LoRA 0.1193/0.0042/0.0042 0.1418/0.0587/0.0587 0.3647/0.4240/0.3927 0.2939/0.3164/0.2753 0.0490/0.0000/0.0000 0.0962/0.0089/0.0089 0.1916/0.2438/0.2320 0.1403/0.0556/0.0556 3
AdaLoRA 0.0079/0.0000/0.0000 0.0848/0.0386/0.0361 0.3548/0.4216/0.3864 0.1814/0.2158/0.1750 0.1443/0.1333/0.1333 0.1408/0.0115/0.0099 0.1828/0.1892/0.1793 0.0957/0.0556/0.0556 1
BayesTune-LoRA 0.0674/0.0075/0.0075 0.0887/0.0420/0.0420 0.3508/0.4174/0.3788 0.2088/0.2442/0.2106 0.1681/0.1905/0.1905 0.1594/0.0089/0.0089 0.1877/0.2362/0.2225 0.0972/0.0556/0.0556 1
FedL2P 0.1093/0.0673/0.0673 0.1171/0.0320/0.0296 0.3895/0.4515/0.4131 0.2629/0.2606/0.2150 0.1349/0.1026/0.1026 0.1335/0.0041/0.0020 0.2011/0.2149/0.2027 0.0645/0.0253/0.0253 1
FedP2EFT 0.1289/0.0165/0.0147 0.1215/0.0166/0.0166 0.4048/0.4910/0.4520 0.2968/0.3065/0.2678 0.2887/0.1667/0.1667 0.0960/0.0069/0.0069 0.2359/0.2863/0.2704 0.0791/0.0222/0.0222 2
@@
Figure 6: Language agnostic rank structure of mBERT in our XNLI setup where the base model is trained with FedIFT full-finetuning ($r=16$).
@@
Figure 7: Language agnostic rank structure of mBERT in our XNLI setup where the base model is trained with FedIFT full-finetuning ($r=2$).
@@
Figure 10: Language agnostic rank structure of mBERT in our MasakhaNEWS setup where the base model is trained with FedIFT full-finetuning ($r=16$).
Figure 11: Language agnostic rank structure of mBERT in our MasakhaNEWS setup where the base model is trained with FedIFT full-finetuning ($r=2$).
Figure 12: Language agnostic rank structure of MobileLLaMA-1.4B in our Fed-Aya setup where the base model is trained with FedIFT LoRA ($r=16$). Zoom in for best results.
Figure 13: Language agnostic rank structure of MobileLLaMA-1.4B in our Fed-Aya setup where the base model is trained with FedIFT LoRA ($r=2$). Zoom in for best results.

### v2

@@ Appendix E Additional Results
$\mathbf{r}$ 2
Lang. LoRA AdaLoRA BT-LoRA FedL2P FedP2EFT
te 0.12/0.05/0.05 0.12/0.06/0.06 0.12/0.06/0.06 0.15/0.07/0.07 0.13/0.06/0.06
ar 0.18/0.02/0.02 0.18/0.02/0.02 0.18/0.02/0.02 0.20/0.02/0.02 0.21/0.03/0.02
es 0.32/0.35/0.33 0.33/0.36/0.33 0.29/0.33/0.31 0.33/0.35/0.32 0.34/0.37/0.34
en 0.25/0.31/0.25 0.25/0.32/0.25 0.23/0.29/0.24 0.26/0.33/0.27 0.27/0.35/0.28
fr 0.20/0.25/0.20 0.20/0.24/0.19 0.19/0.24/0.20 0.23/0.25/0.20 0.23/0.28/0.22
zh 0.06/0.09/0.09 0.07/0.08/0.08 0.06/0.08/0.08 0.05/0.08/0.08 0.05/0.08/0.08
pt 0.26/0.33/0.29 0.27/0.33/0.30 0.24/0.30/0.26 0.28/0.31/0.28 0.28/0.34/0.31
$\mathbf{r}$ 4
Lang. LoRA AdaLoRA BT-LoRA FedL2P FedP2EFT
te 0.12/0.06/0.05 0.12/0.06/0.06 0.12/0.05/0.05 0.13/0.06/0.06 0.14/0.07/0.07
ar 0.19/0.02/0.02 0.19/0.02/0.02 0.19/0.02/0.02 0.18/0.02/0.02 0.22/0.03/0.03
es 0.33/0.36/0.33 0.33/0.35/0.33 0.30/0.34/0.32 0.33/0.36/0.33 0.34/0.37/0.34
en 0.25/0.32/0.26 0.24/0.31/0.25 0.24/0.30/0.24 0.25/0.31/0.25 0.28/0.35/0.29
fr 0.20/0.25/0.20 0.19/0.23/0.19 0.20/0.24/0.19 0.20/0.25/0.20 0.25/0.28/0.22
zh 0.06/0.09/0.09 0.06/0.09/0.09 0.07/0.08/0.08 0.05/0.08/0.08 0.06/0.08/0.08
pt 0.27/0.33/0.30 0.26/0.33/0.29 0.25/0.31/0.27 0.26/0.32/0.28 0.29/0.34/0.31
$\mathbf{r}$ 8
Lang. LoRA AdaLoRA BT-LoRA FedL2P FedP2EFT
te 0.12/0.05/0.05 0.12/0.06/0.06 0.12/0.06/0.06 0.13/0.07/0.07 0.13/0.07/0.06
ar 0.21/0.02/0.02 0.18/0.02/0.02 0.20/0.02/0.02 0.19/0.02/0.02 0.23/0.03/0.02
es 0.34/0.37/0.34 0.32/0.35/0.32 0.32/0.35/0.33 0.33/0.35/0.33 0.35/0.38/0.35
en 0.27/0.34/0.28 0.25/0.31/0.25 0.24/0.31/0.25 0.27/0.33/0.27 0.29/0.37/0.30
fr 0.22/0.26/0.20 0.18/0.23/0.18 0.21/0.25/0.21 0.21/0.26/0.21 0.26/0.28/0.23
zh 0.06/0.09/0.09 0.06/0.08/0.08 0.06/0.08/0.08 0.05/0.08/0.08 0.05/0.08/0.08
pt 0.28/0.34/0.31 0.26/0.32/0.29 0.26/0.32/0.29 0.28/0.33/0.29 0.29/0.35/0.31
$\mathbf{r}$ 16
Lang. LoRA AdaLoRA BT-LoRA FedL2P FedP2EFT
te 0.12/0.06/0.05 0.13/0.06/0.06 0.14/0.07/0.07 0.16/0.08/0.08 0.13/0.07/0.06
ar 0.21/0.02/0.02 0.17/0.02/0.02 0.21/0.02/0.02 0.20/0.02/0.02 0.24/0.03/0.03
es 0.34/0.36/0.34 0.30/0.33/0.30 0.34/0.36/0.34 0.33/0.35/0.33 0.34/0.38/0.35
en 0.28/0.34/0.28 0.24/0.30/0.24 0.26/0.32/0.26 0.28/0.33/0.27 0.28/0.35/0.29
fr 0.25/0.29/0.23 0.18/0.23/0.19 0.22/0.26/0.21 0.23/0.23/0.19 0.25/0.28/0.22
zh 0.06/0.08/0.08 0.06/0.08/0.08 0.05/0.08/0.08 0.05/0.08/0.08 0.05/0.08/0.08
pt 0.29/0.34/0.31 0.25/0.31/0.28 0.29/0.34/0.31 0.28/0.32/0.29 0.30/0.35/0.32
@@
Table 15: Average METEOR/ROUGE-1/ROUGE-L of each language for seen clients of our Fed-Aya setup. The pretrained MobileLLaMA-1.4B model is trained using Standard FL with LoRA following FedLLM-Bench (Ye et al. 2024) and the resulting base model is personalized to each client given a baseline approach.
$\mathbf{r}$ 2
Lang. LoRA AdaLoRA BT-LoRA FedL2P FedP2EFT
te 0.05/0.00/0.00 0.03/0.00/0.00 0.03/0.00/0.00 0.12/0.00/0.00 0.11/0.01/0.01
ar 0.10/0.05/0.05 0.10/0.05/0.05 0.09/0.05/0.05 0.14/0.02/0.02 0.11/0.05/0.05
es 0.35/0.45/0.40 0.32/0.42/0.38 0.36/0.41/0.37 0.39/0.46/0.42 0.38/0.45/0.42
en 0.24/0.27/0.24 0.22/0.26/0.22 0.18/0.20/0.17 0.27/0.30/0.27 0.30/0.34/0.30
fr 0.05/0.00/0.00 0.05/0.05/0.05 0.04/0.00/0.00 0.00/0.00/0.00 0.05/0.00/0.00
zh 0.14/0.01/0.01 0.16/0.01/0.01 0.16/0.01/0.01 0.09/0.01/0.01 0.07/0.01/0.01
pt 0.19/0.22/0.21 0.19/0.23/0.21 0.15/0.16/0.15 0.20/0.22/0.21 0.20/0.25/0.24
ru 0.09/0.06/0.06 0.10/0.06/0.06 0.09/0.06/0.06 0.09/0.06/0.06 0.10/0.06/0.06
$\mathbf{r}$ 4
Lang. LoRA AdaLoRA BT-LoRA FedL2P FedP2EFT
te 0.07/0.01/0.01 0.03/0.00/0.00 0.05/0.01/0.01 0.06/0.01/0.01 0.11/0.02/0.02
ar 0.10/0.06/0.06 0.09/0.06/0.06 0.09/0.05/0.05 0.10/0.06/0.06 0.15/0.03/0.03
es 0.32/0.41/0.37 0.33/0.43/0.39 0.34/0.41/0.36 0.31/0.41/0.38 0.37/0.46/0.43
en 0.23/0.27/0.24 0.18/0.21/0.18 0.17/0.20/0.16 0.25/0.29/0.26 0.29/0.33/0.28
fr 0.05/0.05/0.05 0.05/0.00/0.00 0.12/0.06/0.06 0.12/0.07/0.07 0.05/0.00/0.00
zh 0.15/0.01/0.01 0.16/0.01/0.00 0.16/0.01/0.01 0.16/0.00/0.00 0.07/0.01/0.01
pt 0.19/0.23/0.21 0.18/0.21/0.20 0.15/0.16/0.15 0.17/0.19/0.18 0.19/0.25/0.24
ru 0.10/0.06/0.06 0.12/0.06/0.06 0.09/0.02/0.02 0.08/0.06/0.06 0.13/0.06/0.06
$\mathbf{r}$ 8
Lang. LoRA AdaLoRA BT-LoRA FedL2P FedP2EFT
te 0.10/0.00/0.00 0.02/0.00/0.00 0.05/0.00/0.00 0.09/0.00/0.00 0.11/0.01/0.01
ar 0.12/0.05/0.05 0.09/0.04/0.04 0.10/0.06/0.06 0.11/0.05/0.05 0.16/0.01/0.01
es 0.34/0.44/0.40 0.34/0.43/0.39 0.35/0.43/0.38 0.32/0.44/0.41 0.45/0.52/0.49
en 0.27/0.31/0.27 0.18/0.21/0.18 0.19/0.22/0.19 0.24/0.28/0.25 0.28/0.29/0.25
fr 0.05/0.00/0.00 0.12/0.08/0.08 0.13/0.08/0.08 0.05/0.00/0.00 0.05/0.00/0.00
zh 0.12/0.01/0.01 0.15/0.01/0.00 0.15/0.01/0.01 0.14/0.01/0.01 0.11/0.01/0.01
pt 0.19/0.23/0.22 0.18/0.19/0.18 0.14/0.16/0.15 0.19/0.23/0.21 0.22/0.27/0.26
ru 0.08/0.06/0.06 0.11/0.06/0.06 0.09/0.06/0.06 0.12/0.06/0.06 0.09/0.05/0.05
$\mathbf{r}$ 16
Lang. LoRA AdaLoRA BT-LoRA FedL2P FedP2EFT
te 0.12/0.00/0.00 0.01/0.00/0.00 0.07/0.01/0.01 0.11/0.07/0.07 0.13/0.02/0.01
ar 0.14/0.06/0.06 0.08/0.04/0.04 0.09/0.04/0.04 0.12/0.03/0.03 0.12/0.02/0.02
es 0.36/0.42/0.39 0.35/0.42/0.39 0.35/0.42/0.38 0.39/0.45/0.41 0.40/0.49/0.45
en 0.29/0.32/0.28 0.18/0.22/0.17 0.21/0.24/0.21 0.26/0.26/0.21 0.30/0.31/0.27
fr 0.05/0.00/0.00 0.14/0.13/0.13 0.17/0.19/0.19 0.13/0.10/0.10 0.29/0.17/0.17
zh 0.10/0.01/0.01 0.14/0.01/0.01 0.16/0.01/0.01 0.13/0.00/0.00 0.10/0.01/0.01
pt 0.19/0.24/0.23 0.18/0.19/0.18 0.19/0.24/0.22 0.20/0.21/0.20 0.24/0.29/0.27
ru 0.14/0.06/0.06 0.10/0.06/0.06 0.10/0.06/0.06 0.06/0.03/0.03 0.08/0.02/0.02
@@
Table 16: Average METEOR/ROUGE-1/ROUGE-L of each language for unseen clients of our Fed-Aya setup. The pretrained MobileLLaMA-1.4B model is trained using Standard FL with LoRA following FedLLM-Bench (Ye et al. 2024) and the resulting base model is personalized to each client given a baseline approach.
@@
Figure 5: Language agnostic rank structure of mBERT in our XNLI setup where the base model is trained with Standard FL full-finetuning ($r=16$).
@@
Figure 6: Language agnostic rank structure of mBERT in our XNLI setup where the base model is trained with Standard FL full-finetuning ($r=2$).
@@
Figure 9: Language agnostic rank structure of mBERT in our MasakhaNEWS setup where the base model is trained with Standard FL full-finetuning ($r=16$).
Figure 10: Language agnostic rank structure of mBERT in our MasakhaNEWS setup where the base model is trained with Standard FL full-finetuning ($r=2$).
Figure 11: Language agnostic rank structure of MobileLLaMA-1.4B in our Fed-Aya setup where the base model is trained with Standard FL LoRA ($r=16$). Zoom in for best results.
Figure 12: Language agnostic rank structure of MobileLLaMA-1.4B in our Fed-Aya setup where the base model is trained with Standard FL LoRA ($r=2$). Zoom in for best results.
@@
Figure 15: Cross-lingual $\bm{\lambda}$ distance among languages in our MasakhaNEWS setup. Each block shows the log-scale normalized average Euclidean distances between all pairs of clients’ $\bm{\lambda}$ in their respective languages. The smaller the distance, the more similar $\bm{\lambda}$ is.